# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 2.4 Generic Coordinates
Theorem 2.5 ([CG21, Theorem 9, 10, Corollary 2]).
In the Gröbner basis and algebraic cryptanalysis literature the latter inequality is also usually referred as Macaulay bound. Moreover, Lazard [Laz83] first observed that DRL Gröbner bases for “generic” polynomial systems can be found within the Macaulay bound.
Finally, being in generic coordinates was recently characterized by Steiner [Ste24].
Theorem 2.6 ([Ste24, Theorem 3.2]).
Let $K$ be an algebraically closed field, and let =$\mathcal{F}=\{f_{1},\dots,f_{m}\}\subset K[x_{1},\dots,x_{n}]$ be a polynomial system such that
  1. (i)
    $(\mathcal{F})\neq(1)$, and
  2. (ii)
    $\dim\left(\mathcal{F}\right)=0$.
Then the following are equivalent.
  1. (1)
    $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates.
  2. (2)
    $\sqrt{\mathcal{F}^{\text{top}}}=\left(x_{1},\dots,x_{n}\right)$.
  3. (3)
    $\left(\mathcal{F}^{\text{top}}\right)$ is zero-dimensional in $K[x_{1},\dots,x_{n}]$.
  4. (4)
    For every $1\leq i\leq n$ there exists $d_{i}\in\mathbb{Z}_{\geq 1}$ such that $x_{i}^{d_{i}}\in\operatorname{in}_{DRL}\left(\mathcal{F}^{\text{hom}}\right)$.
Let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system, to prove that $\mathcal{F}^{\text{hom}}$ is in generic coordinates one can use the following procedure based on (2):
@@
Another quantity in the literature that is often used to express the complexity of finding a DRL Gröbner basis is the so-called degree of regularity.
Definition 2.7 (Degree of regularity, [BFS04, Definition 4]).
@@
Let $K$ be a field, and let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$. Assume that $\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}$ for some integer $d\geq 0$. The degree of regularity is defined as
$d_{\operatorname{reg}}\left(\mathcal{F}\right)=\min\left\{d\geq 0\;\middle|\;\left(\mathcal{F}^{\text{top}}\right)_{d}=P_{d}\right\}.$
Note that by Theorem 2.6 a polynomial system is in generic coordinates if and only if it admits a finite degree of regularity.
@@ 2.5 The Eigenvalue Method
As outlined in the introduction, in the algebraic cryptanalysis literature the standard approach to polynomial system solving is
  1. (1)
    DRL Gröbner basis computation,
  2. (2)
    term order conversion to LEX, and
  3. (3)
    univariate factoring.
In this section we shortly discuss that there is an alternative to term order conversion via linear algebra-based techniques [KR16, Chapter 6]. Our motivation to change the strategy is quite simple. Via the Eigenvalue Method we will be able to exploit the multiplicative structure of Ciminion and Hydra DRL Gröbner bases. Consequently, this will provide a slight improvement compared to the generic complexity of the probabilistic FGLM algorithm [FGHR14].
Let $I\subset P=K[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, i.e. $P/I$ is a finite dimensional $K$-vector space [KR00, Proposition 3.7.1], and let $d=\dim_{K}\left(P/I\right)$. If $>$ is a term order on $P$ and $\mathcal{G}\subset I$ is $>$-Gröbner basis, then a $K$-vector space basis $\mathcal{B}$ of $P/I$ is given by the monomials not contained in $\operatorname{in}_{>}\left(\mathcal{G}\right)=\operatorname{in}_{>}\left(I\right)$. Now let $f\in P$, then the multiplication map
$\begin{split}\theta_{f}:P&\to P/I,\\ x&\mapsto x\cdot f,\end{split}$ (23)
can be represented via a matrix. This matrix is called the multiplication matrix $\mathbf{M}_{f}$ for $f$. Given a $>$-Gröbner basis its computation is straight-forward: Index the columns of $\mathbf{M}_{f}$ by the $K$-vector space basis $\mathcal{B}$, and the rows by $b\cdot f$, where $b\in\mathcal{B}$. The row $b\cdot f$ is then simply the coefficient vector of $b\cdot f$ in $P/I$ which can be computed via $b\cdot f\mod(\mathcal{G})$ with respect to $>$.
Now assume in addition that $K$ is algebraically closed, then for every point $\mathbf{x}\in\mathcal{Z}\left(I\right)$ the point $f(\mathbf{x})$ is an eigenvalue of $\mathbf{M}_{f}$, see [KR16, Proposition 6.2.1]. Therefore, we can compute the zero locus $\mathcal{Z}\left(I\right)$ as follows: Compute the eigenvalues of $\mathbf{M}_{x_{1}},\dots,\mathbf{M}_{x_{n}}$, iterate over all possible combinations, and verify whether a combination is indeed a solution to $I$. This approach is known as Eigenvalue Method [KR16, §6.2.A].
In his seminal work Strassen showed that matrix inversion, determinant and Gaussian elimination have the same complexity as matrix multiplication up to some constant [Str69]. For the determinant this observation follows form LU-decomposition of block square matrices. We recall this decomposition since we will exploit it at a later point in this paper. Let $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}\in K^{2\cdot N\times 2\cdot N}$ be matrices, then we have the block matrix decomposition
$\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\begin{pmatrix}\mathbf{I}_{N\times N}&\mathbf{0}_{N\times N}\\ \mathbf{C}\mathbf{A}^{-1}&\mathbf{I}_{N\times N}\end{pmatrix}\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{0}_{N\times N}&\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\end{pmatrix},$ (24)
provided that $\mathbf{A}$ is invertible. With well-known formulae for block matrix determinants, see e.g. [Sil00], one then has that
$\det\begin{pmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{pmatrix}=\det\left(\mathbf{A}\right)\cdot\det\left(\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}\right).$ (25)
So the determinant of the block matrix has been reduced to two determinants of smaller matrices, one matrix inversion and two matrix multiplications. By Strassen [Str69] the complexity of multiplying two $N\times N$ matrices is in $\mathcal{O}\left(N^{\omega+o(1)}\right)$, where $2.372862\leq\omega<$ 2.37286 is again the linear algebra constant [AW21]. Note that the term $N^{o(1)}$ is usually ignored in complexity estimations.
Lastly, it is well known that the eigenvalues of a matrix $\mathbf{M}\in K^{N\times N}$ are roots of its characteristic polynomial
@@
Therefore, construction of the characteristic polynomials in the Eigenvalue Method can be performed in

### v4

@@ 2.3.2 Polynomial Systems in Generic Coordinates
Let $\mathcal{F}\subset K[x_{1},\dots,x_{n}]$ be an inhomogeneous polynomial system, to prove that $\left(\mathcal{F}^{\text{hom}}\right)$ is in generic coordinates one can use the following procedure based on [Ste24, Theorem 3.2]:
@@
Theorem 2.6 ([CG21, Theorem 9, 10, Corollary 2]).
In the Gröbner basis and algebraic cryptanalysis literature, the latter inequality is often referred to as the Macaulay bound.
2.3.3 Complexity Estimates via the Solving Degree
It is well-known that the complexity of Gaussian elimination of a matrix $\mathbf{A}\in K^{n\times m}$ with $r=\operatorname{rank}\left(\mathbf{A}\right)$ is given by $\mathcal{O}\left(n\cdot m\cdot r^{\omega-2}\right)$ field operations [Sto00, §2.2], where $2.3715522\leq\omega<$ 2.371552 is a linear algebra constant [WXXZ24].
Given an inhomogeneous polynomial system $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$, it is well-known that the number of monomials in $P$ of degree $d$ is $\binom{n+d-1}{d}$. Thus, the number of columns of $M_{\leq d}$ is bounded by $\sum_{i=0}^{d}\binom{n+i-1}{i}=\binom{n+d}{d}$, and the number of rows by $\left|\mathcal{F}\right|\cdot\binom{n+d}{d}$. So, the complexity of Gaussian elimination on the inhomogeneous Macaulay matrices $M_{\leq d}$ is bounded by
$\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\binom{n+d}{d}^{\omega}\right).$ (14)
2.4 Boolean Macaulay Matrices
Let $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be a polynomial system. If we are only interested in $\mathbb{F}_{2}$-valued solutions, then it is sufficient to compute a Gröbner basis of the ideal $\big{(}\mathcal{F},x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\big{)}$. On the other hand, via the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ we can reduce any monomial in the Macaulay matrix to a square-free monomial. Effectively, this reduces the full Macaulay matrix to a square-free submatrix, the so-called Boolean Macaulay matrix. Gaussian elimination for Boolean Macaulay matrices with respect to DRL was analyzed in [BFSS13] by Bardet et al.
Interestingly, for Ciminion and Hydra we will construct zero-dimensional DRL Gröbner bases in Sections 3.1 and 4.2 with leading monomials $x_{i}^{2}$. Hence, for Gröbner basis computations we can use these bases to reduce the Macaulay matrix to a “Boolean” one over an arbitrary finite field $\mathbb{F}_{q}$. This motivates the following definitions.
Definition 2.7.
  1. (i)
  2. (ii)
    $\mathcal{F}$ is a $>$-Gröbner basis.
Note that Condition (ii) is introduced so that the remainder of multivariate polynomial division does not depend on the ordering of the elements of $\mathcal{F}$, see [CLO15, Chapter 2 §6 Proposition 1].
Definition 2.8 (cf. [BFSS13, Definition 1]).
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. For the (inhomogeneous) $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$:
  • Columns are indexed by square-free monomials $s\in P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(m\right)\leq d$ sorted via $>$.
  • Rows are indexed by polynomials $t\cdot f$ where $t\in P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ is a square-free monomial and $f\in\mathcal{F}$ such that $\deg\left(t\cdot f\right)\leq d$.
Note that any monomial $t\in P$ can be reduced to a square-free polynomial $\hat{t}\equiv t\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(\hat{t}\right)\leq\deg\left(t\right)$ due to degree compatibility of $>$. Since remainders with respect to a Gröbner basis are unique, see [CLO15, Chapter 2 §6 Proposition 1], for the remainder of $t\cdot f$ it does not matter whether we reduce $t$ first and then $\hat{t}\cdot f$ or $t\cdot f$ at once. So it is indeed sufficient to consider only products with square-free monomials.
Obviously, the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ is a polynomial consisting of only square-free terms. Moreover, by elementary properties of multivariate polynomial division with respect to a degree compatible term order, see [CLO15, Chapter 2 §3 Theorem 3], we have that
$t\cdot f=\sum_{i=1}^{m}g_{i}\cdot f_{i}^{\text{Bool}}+r_{t\cdot f},$ (15)
where $\deg\left(g_{i}\cdot f_{i}^{\text{Bool}}\right),\deg\left(r_{t\cdot f}\right)\leq\deg\left(t\cdot f\right)$. I.e., the $K$-Boolean Macaulay matrix can be constructed via row operations from the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$. Alternatively, we can obtain the $K$-Boolean Macaulay matrix by pivoting non-square-free monomials in $M_{\leq d}$ via rows coming from $\mathcal{F}_{\text{Bool}}$, and then performing an auxiliary Gaussian elimination step that stops after eliminating all non-square-free monomials in the rows coming from $\mathcal{F}$.
Let $M_{\leq d}$ be a Macaulay matrix, we denote with $\operatorname{rowsp}\left(M_{\leq d}\right)$ the row space basis, i.e. the polynomial system obtained via Gaussian elimination on $M_{\leq d}$. Analogously, we use this notation for Boolean Macaulay matrices.
Definition 2.9.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,\allowbreak x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. The $K$-Boolean solving degree of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ is the least degree $d$ such that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ contains a Gröbner basis with respect to $>$. We denote it by $\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
Note that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ might not contain a zero-dimensional Gröbner basis of the full system. A trivial example is =$\mathcal{F}=\left\{x_{1}^{2},x_{1}\cdot x_{2},x_{2}^{2}\right\}$. Therefore, to extract the Gröbner basis from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ we set up the ideal of $>$-leading terms and then compute its minimal generating set. With the minimal generating set we can then filter the $>$-Gröbner basis.444 It is obvious that all leading terms of $\operatorname{rowsp}\left(M_{\leq d}\right)$ not coming from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ are already contained in $\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$. Since $\mathcal{F}_{\text{Bool}}$ is a $>$-Gröbner basis it is indeed sufficient to just consider $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$.
Theorem 2.10.
@@ 2.4 Boolean Macaulay Matrices
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Let $D=\min_{f\in\mathcal{F}}\deg\left(f\right)$, and let $d=\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. Then
  1. (1)
  2. (2)
    If $d-D<n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires
    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\left(\sum_{i=0}^{d-D}\binom{n}{i}\right)\cdot\left(\sum_{i=0}^{d}\binom{n}{i}\right)^{\omega-1}\right)$
    field operations.
  3. (3)
    If $d-D\geq n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires at most
    field operations.
Proof.
For (1), in $M_{\leq d}$ we pivot non-square-free columns via rows coming from $\mathcal{F}_{\text{Bool}}$, then we perform the auxiliary Gaussian elimination step that produces $M_{\leq d}^{\text{Bool}}$ as submatrix, and finally we perform Gaussian elimination on $M_{\leq d}^{\text{Bool}}$. Since Gaussian elimination does not depend on the order of reductions $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ has to contain the $>$-Gröbner basis.
For (2), it is well-known that in degree $i$ there exist $\binom{n}{i}$ square-free monomials in $P$. The polynomial $f\in\mathcal{F}$ contributes $\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}$ many rows to the $K$-Boolean Macaulay matrix. Then, the total number of rows in $M_{\leq d}^{\text{Bool}}$ is bounded by
$\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}\leq\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-D}\binom{n}{i}=\left|F\right|\cdot\sum_{i=0}^{d-D}\binom{n}{i}.$
The number of columns is bounded by $\sum_{i=0}^{d}\binom{n}{i}$, so we also have that $\operatorname{rank}\left(M_{\leq d}^{\text{Bool}}\right)\leq\sum_{i=0}^{d}\binom{n}{i}$, and the claim follows from the complexity of Gaussian elimination [Sto00, §2.2].
For (3), if $d-D\geq n$, then there exists at least one polynomial $f\in\mathcal{F}$ which has to be multiplied with any square-free monomial $s\in P$. So, $f$ contributes $\sum_{i=0}^{n}\binom{n}{i}=2^{n}$ many rows. Obviously, any other $g\in\mathcal{F}$ contributes at most $\leq 2^{n}$ many rows. Hence, the total number of rows is bounded by $\left|\mathcal{F}\right|\cdot 2^{n}$, and the number of columns as well as the rank are bounded by $2^{n}$. Again, the claim follows by [Sto00, §2.2]. ∎
Remark 2.11.
This theorem can be refined by replacing the binomial coefficients $\binom{n}{i}$ with the number of monomials in $P\setminus\operatorname{LM}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ of degree $i$.
So far our analysis does not differ from the one over $\mathbb{F}_{2}$, however the binary field has a special feature. Division with respect to the field equations $\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ is “for free”, i.e. its complexity is negligible compared to Gaussian elimination. For general $K$-Boolean polynomial systems this is not necessarily the case. The complexity of general polynomial division with respect to quadratic $K$-Boolean polynomial systems follows as corollary to [Ten19, Proposition 3.10].
Corollary 2.12.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}}\subset P=K[x_{1},\dots,x_{n}]$ be a quadratic polynomial system, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$.
  1. (1)
    Let $f\in P$, then computation of $r_{f}=f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$ requires
    $\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\binom{n+\deg\left(f\right)}{\deg\left(f\right)}\right)$
    field operations.
  2. (2)
    Let $\mathcal{F}\subset P$, $D_{\min}=\min_{f\in\mathcal{F}}\deg\left(f\right)$ and $D_{\max}=\max_{f\in\mathcal{F}}\deg\left(f\right)$, then construction of $M_{\leq d}^{\text{Bool}}$ for $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ requires
    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i}\right)$
    field operations.
Proof.
For (1), in $P$ there exist $\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}=\binom{n+\deg\left(f\right)}{\deg\left(f\right)}$ monomials of degree $\leq\deg\left(f\right)$ and $\binom{n+1}{2}+\binom{n}{1}+1$ monomials of degree $\leq 2$. Now the claim follows from [Ten19, Proposition 3.10].
For (2), with (1) we can estimate the complexity of the construction as
$\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-\deg\left(f\right)}\binom{n}{i}\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\binom{n+\deg\left(f\right)+i}{\deg\left(f\right)+i}$
$\stackrel{{\scriptstyle(\ast)}}{{\leq}}\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{f\in\mathcal{F}}\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i}$
$=\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{d-D_{\min}}\binom{n}{i}\cdot\binom{n+D_{\max}+i}{D_{\max}+i},$
where ()($\ast$)( ∗ ) follows from $d-\deg\left(f\right)\leq d-D_{\min}$ and the fact that $\binom{n+d}{d}$ is strictly increasing in $d$, hence $\binom{n+\deg\left(f\right)+i}{\deg\left(f\right)+i}\leq\binom{n+D_{\max}+i}{D_{\max}+i}$. ∎
Of course, the $K$-Boolean Macaulay matrix generalizes to more advanced linear algebra-based Gröbner basis algorithms like F4/5 [Fau99, Fau02].
@@ 2.5 The Eigenvalue Method
Now we discuss how the zeros of a polynomial system can be found via linear algebra once a zero-dimensional DRL Gröbner basis of the input system is known [KR16, Chapter 6]. Let $I\subset P=K[x_{1},\dots,x_{n}]$ be a zero-dimensional ideal, i.e. $P/I$ is a finite dimensional $K$-vector space [KR00, Proposition 3.7.1], and let $d=\dim_{K}\left(P/I\right)$. If $>$ is a term order on $P$ and $\mathcal{G}\subset I$ is $>$-Gröbner basis, then a $K$-vector space basis $\mathcal{B}$ of $P/I$ is given by the monomials not contained in $\operatorname{LM}_{>}\left(\mathcal{G}\right)=\operatorname{LM}_{>}\left(I\right)$. Now let $f\in P$, then the multiplication map
$\theta_{f}:P/I\to P/I,\qquad\qquad x\mapsto x\cdot f,$ (16)
is $K$-linear, so it can be represented by a matrix. This matrix is called the multiplication matrix $\mathbf{M}_{f}$ for $f$. Given a $>$-Gröbner basis its computation is straight-forward: Index the columns of $\mathbf{M}_{f}$ by the $K$-vector space basis $\mathcal{B}$, and the rows by $b\cdot f$, where $b\in\mathcal{B}$. The row $b\cdot f$ is then simply the coefficient vector of $b\cdot f$ in $P/I$, which can be computed via $b\cdot f\mod(\mathcal{G})$ with respect to $>$.
Now assume in addition that $K$ is algebraically closed, then for every point in the variety $\mathbf{x}\in\mathcal{V}\left(I\right)=\left\{\mathbf{x}\in K^{n}\mid\forall g\in I\!:g(\mathbf{x})=0\right\}$ the point $f(\mathbf{x})$ is an eigenvalue of $\mathbf{M}_{f}$, see [KR16, Proposition 6.2.1]. Therefore, we can compute the variety $\mathcal{V}\left(I\right)$ as follows: Compute the eigenvalues of $\mathbf{M}_{x_{1}},\dots,\mathbf{M}_{x_{n}}$, iterate over all possible combinations, and verify whether a combination is indeed a solution to $I$. This approach is known as Eigenvalue Method [KR16, §6.2.A].
It is well-known that the eigenvalues of a matrix $\mathbf{M}\in K^{N\times N}$ are given by the roots of its characteristic polynomial
@@
So the complexity of the Eigenvalue Method boils down to $n$ determinant computations with polynomial entries.
2.5.1 Solving Structured Polynomial Systems Over Finite Fields
Let =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset P=K[x_{1},\dots,x_{n}]$ be such that $\operatorname{LM}_{DRL}\left(f_{i}\right)=x_{i}^{d_{i}}$ for all $1\leq i\leq n$. Then, trivially $\mathcal{F}$ is a DRL Gröbner basis, see e.g. [CLO15, Chapter 2 §9 Theorem 3, Proposition 4] with $\dim_{K}\left(\mathcal{F}\right)=\prod_{i=1}^{n}d_{i}$. Let $D=\prod_{i=1}^{n-1}d_{i}$, and let $\mathcal{B}^{\prime}\subset K[x_{1},\dots,x_{n-1}]/\left(x_{1}^{d_{1}},\dots,x_{n-1}^{d_{n-1}}\right)$ be some $K$-vector space basis, then $\mathcal{B}=\bigcup_{i=0}^{d_{n}-1}x_{n}^{i}\cdot\mathcal{B}^{\prime}$ is a $K$-vector space basis for $P/(\mathcal{F})$. In particular, the multiplication matrix $\mathbf{M}_{x_{n}}$ is then of the form
[…]
where $\mathbf{0}\in K^{D\times D}$ denotes the zero matrix, $\mathbf{I}\in K^{D\times D}$ the identity matrix, and $\mathbf{A}_{0},\dots,\allowbreak\mathbf{A}_{d_{n}-1}\in K^{D\times D}$ some matrices. By [BBL+24, Lemma 2] we then have that
$\chi_{\mathbf{M}_{x_{n}}}(x)=\det\left(x_{n}^{d_{n}}\cdot\mathbf{I}-\sum_{i=0}^{d_{n}-1}x_{n}^{i}\cdot\mathbf{A}_{i}\right),$ (19)
and computation of this determinant requires [BBL+24, §3.2]
$\mathcal{O}\bigg{(}d_{n}\cdot\log_{2}(d_{n})^{2}\cdot\Big{(}1+\log_{2}\big{(}\log_{2}(d_{n})\big{)}\Big{)}\cdot\Big{(}\frac{D}{d_{n}}\Big{)}^{\omega}\bigg{)}$ (20)
field operations, where $2.3715522\leq\omega<$ 2.371552 is again the linear algebra constant [WXXZ24].
In this paper we solely work over finite fields $\mathbb{F}_{q}$, and we are only interested in the $\mathbb{F}_{q}$-valued roots of $\mathcal{F}$. Given a univariate polynomial $f\in\mathbb{F}_{q}$ with $\deg\left(f\right)=d$ we can always achieve this restriction by passing to the GCD with the field equation $g=\gcd\left(f,x^{q}-x\right)$. By [BBLP22, §3.1] computation of this GCD requires
$\mathcal{O}\left(\begin{rcases}\begin{dcases}d\cdot\log_{2}(d)\cdot\log\big{(}\log(d)\big{)}\cdot\big{(}\log(d)+\log_{2}(q)\big{)},&d<q,\\ d\cdot\log(d)^{2}\cdot\log\big{(}\log(d)\big{)},&d\geq q\end{dcases}\end{rcases}\right)$ (21)
field operations. Provided that only a few roots of $f$ come from $\mathbb{F}_{q}$ the final factoring step can be neglected.
In this paper we will encounter quadratic DRL Gröbner bases =$\mathcal{F}=\{f_{1},\dots,f_{n}\}\subset P=\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ of the form
$f_{i}=\mathcal{Q}\left(x_{i},\dots,x_{n}\right)+\mathcal{A}_{i}\left(x_{1},\dots,x_{n}\right),$ (22)
where $\mathcal{Q}_{i}\in P$ is a quadratic polynomial which has the term $x_{i}^{2}$ and $\mathcal{A}_{i}\in P$ is an affine polynomial which has the term $x_{n-i+1}$ for $1\leq i\leq n$. It is easy to see that $\mathcal{F}$ is a DRL Gröbner basis for $x_{1}>\ldots>x_{n}$. For such systems we can give an optimized Eigenvalue Method to find the $\mathbb{F}_{q}$-valued solutions. We use a $\mathbb{F}_{q}$-vector space basis such that the multiplication matrix $\mathbf{M}_{x_{n}}$ has the shape from Equation 18. Then, the characteristic polynomial is constructed in $\mathcal{O}\left(2^{\omega\cdot(n-1)+1}\right)$ field operations, and its $\mathbb{F}_{q}$-valued roots are extracted via the GCD with the field equation. Let $\alpha\in\mathbb{F}_{q}$ be such a solution for $x_{n}$, then we can substitute $x_{n}=\alpha$ back into $\mathcal{F}$. In particular, from $f_{n}$ we obtain the affine equation
$\alpha^{2}+\tilde{\mathcal{A}_{n}}=0.$ (23)
Now we rearrange this equation to have $x_{1}$ on the right-hand side. Then, we can additionally eliminate the variable $x_{1}$ in the equations $f_{1}=\ldots=f_{n-1}=0$. Since $x_{1}$ is not present in the quadratic terms of $f_{2},\dots,f_{n-1}\in\mathbb{F}_{q}[x_{2},\dots,x_{n-1}]$, they are again a zero-dimensional DRL Gröbner basis. The polynomial $f_{1}$ is an additional overdetermined polynomial which we ignore. Now we iterate this procedure to solve for $x_{n-1},x_{n-2},\dots,x_{\left\lfloor{\frac{n}{2}}\right\rfloor}$, then we lift to a global solution guess for $x_{1},\dots,x_{n}$, and finally verify it on $\mathcal{F}$. Let $\mathcal{D}(d,q)$ denote the complexity of the GCD between a degree $d$ polynomial and the field equation of $\mathbb{F}_{q}$, moreover let $N$ be a bound on the number of $\mathbb{F}_{q}$-valued solutions at each iteration in our solving algorithm. Then, this dedicated Eigenvalue Method requires